OpenStudy (anonymous):

A refrigerator with a weight of 1,127 newtons needs to be moved into a house using a ramp. The length of the ramp is 2.1 meters, and its height is 0.85 meters. What is the percentage efficiency if a force of 496 newtons is applied to the refrigerator?

OpenStudy (anonymous):

okay this one i got 85% but it didnt match my options, any opinions?

OpenStudy (anonymous):

I think it should me 91%. What is the correct ans?

OpenStudy (anonymous):

Well one of my options is 92%? How did you solve it?

OpenStudy (anonymous):

Yes, 92%. What was the work done by agent. He applied force 496N ALONG the ramp. Since ramp is 2.1m long So work done by him =2.1*496 = 1041.6 J gain in P.E =weight*height increase =1127*0.85 =957 J Divide and find the efficiency.

OpenStudy (anonymous):

Ohh okay, i saw my mistake. Thanks for the help :)